With sickle cell anemia, being heterozygous gives an advantage against malaria infection in endemic areas. In a part of Tanzania, the relative finesses of AA, AS and SS genotypes have been estimated to be 0.68, 1 and 0.37.

Calculate the equilibrium frequency of q^ of the sickle allele
Answer #1

Here relative inesses of AA = 0.68

for AS = 1

for SS = 0.37

so here it is given that

p2 : q2 ; ;  0.68 ; 0.37

taking square root

p : q ; ; 0.8246 : 0.6083

so that means

p/q = 0.8246/0.6083

p/q = 1.3556

p = 1.3556q

so here as we know that

p + q = 1

1.3556 q + q = 1

2.3556 q = 1

q = 0.4245

so the equilibrim frequency of q^ of the sickle allele = 0.4245